OpenStudy (anonymous):

its 7/5/9 subtract 3/7/9 will give medals

OpenStudy (yolo_king):

Do you mean \[7\frac{ 5 }{ 9 } -3\frac{ 7 }{ 9 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know how to solve it

OpenStudy (yolo_king):

What do you think is the answer?

OpenStudy (yolo_king):

OR before you can add or subtract fractions with different denominators, you must first find equivalent fractions with the same denominator, like this: 1.Find the smallest multiple (LCM) of both numbers. 2.Rewrite the fractions as equivalent fractions with the LCM as the denominator.

OpenStudy (anonymous):

I put it was 3/7/9

OpenStudy (yolo_king):

Th\[\frac{ 34 }{ 9 }\]e answer is

OpenStudy (anonymous):

I don't under stand

OpenStudy (eric_d):

Welcome to Open Study @Raygirl18

OpenStudy (anonymous):

thank you

OpenStudy (eric_d):

Change the given fractions which are in mixed number form into improper fraction @Raygirl18

OpenStudy (eric_d):



OpenStudy (anonymous):

I don't know how to do that im really bad at math

OpenStudy (eric_d):

Don't worry @Raygirl18 I'll teach you..

OpenStudy (anonymous):

aww thank you

OpenStudy (eric_d):